OpenStudy (anonymous):

Let Log p/n = 6 Log m/n = 8 What is the relationship between P and M? A. M=100P B. M= 2P C. M=.001P D. P=100M

OpenStudy (anonymous):

logp-logn=6 logm-logn=8 now subtract 1 from 2

OpenStudy (anonymous):

logp-logn=6...............1 logm-logn=8...............2 now subtract 1 from 2

OpenStudy (anonymous):

I'm still a bit confused

OpenStudy (anonymous):

Oh so would it be a? M=100 p

OpenStudy (anonymous):

logm-log p=8-6=2 log m/p=2 10^2=m/p m=100p yes